OpenStudy (anonymous):

is 3 - 2/ 3+1/4 all over 3/4 - 2/ 1-1/3 = -1 ,5/39? thankyou so much guyss,, i really learned a lot here ^^ thanks

OpenStudy (anonymous):

cant understand what the question is

OpenStudy (anonymous):

\[\frac{ 3 - \frac{ 2 }{ 3 } + \frac{ 1 }{ 4 } }{ \frac{ 3 }{ 4 }- \frac{ 2 }{ 1 }-\frac{ 1 }{ 3 } } \]Is this your question?

OpenStudy (anonymous):

 this is the question @miteshchvm and @curiousshubham :) thanks